OpenStudy (anonymous):

-1/3. (8 1/7) simplify

OpenStudy (anonymous):

\[\frac{ -1 }{ 3 }*\frac{ 57 }{ 7 }=\frac{ -57 }{ 21 }= -2 \frac{ 15 }{ 21 } \]

OpenStudy (anonymous):

-19/7 = -2 5/7

OpenStudy (anonymous):

thank you